### v4

@@ Abstract
The rapid advancement of large models, driven by their exceptional abilities in learning and generalization through large-scale pre-training, has reshaped the landscape of Artificial Intelligence (AI). These models are now foundational to a wide range of applications, including conversational AI, recommendation systems, autonomous driving, content generation, medical diagnostics, and scientific discovery. However, their widespread deployment also exposes them to significant safety risks, raising concerns about robustness, reliability, and ethical implications. This survey provides a systematic review of current safety research on large models, covering Vision Foundation Models (VFMs), Large Language Models (LLMs), Vision-Language Pre-training (VLP) models, Vision-Language Models (VLMs), Diffusion Models (DMs), and large-model-based Agents. Our contributions are summarized as follows: (1) We present a comprehensive taxonomy of safety threats to these models, including adversarial attacks, data poisoning, backdoor attacks, jailbreak and prompt injection attacks, energy-latency attacks, data and model extraction attacks, and emerging agent-specific threats. (2) We review defense strategies proposed for each type of attacks if available and summarize the commonly used datasets and benchmarks for safety research. (3) Building on this, we identify and discuss the open challenges in large model safety, emphasizing the need for comprehensive safety evaluations, scalable and effective defense mechanisms, and sustainable data practices. More importantly, we highlight the necessity of collective efforts from the research community and international collaboration. Our work can serve as a useful reference for researchers and practitioners, fostering the ongoing development of comprehensive defense systems and platforms to safeguard AI models. GitHub: https://github.com/xingjunm/Awesome-Large-Model-Safety.
@@
Large Model Safety, AI Safety, Attacks and Defenses
@@ 1 Introduction
Figure 1: Left: The number of safety research papers published over the past four years. Middle: The distribution of research across different models. Right: The distribution of research across different types of attacks and defenses.
Figure 2: Left: The quarterly trend in the number of safety research papers published across different models; Middle: The proportional relationship between large models and their corresponding attacks and defenses; Right: The annual trend in the number of safety research papers published on various attacks and defenses, presented in descending order from highest to lowest.
@@
forked edges, for tree= grow=east, reversed=true, anchor=mid west, parent anchor=mid east, child anchor=mid west, base=left, font=, rectangle, draw=none, sharp corners, align=left, minimum width=2em, edge+=darkgray, line width=0.6pt, s sep=2.8pt, inner xsep=1.8pt, inner ysep=2.7pt, line width=0.7pt, par/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=0font=,, where level=1text width=3.8em,font=,, where level=2text width=8.5em,font=,, [ Large Model Safety, par [ Vision Foundation Models (§ 2) , vfm-bg-node [ Attacks and Defenses for ViT (§ 2.1), vfm-line-node [ (i) Adversarial Attacks: [1] [2] [3] [4] [5] [6] [7] [8] [9] [10] [11] [12] [13] [14] [15] [16] [17] [18] [19]
@@
(iii) Backdoor Attacks: [39] [40] [41] [42] [43](iv) Backdoor Defenses: [44] [45] [46] , vfm-line-node, text width=27.9em ] ] [ Attacks and Defenses for SAM (§ 2.2), vfm-line-node [ (i) Adversarial Attacks: [47] [48] [49] [50] [51] [52] [53] [54](ii) Adversarial Defenses: [55]
@@
(iii) Backdoor & Poisoning Attacks: [56] [57] , vfm-line-node, text width=27.9em ] ] ] [ Large Language Models (§ 3) , llm-bg-node [ Adversarial Attack (§ 3.1), llm-line-node [ (i) White-Box: [58] [59] [60] [61] [62] [63] [64](ii) Black-Box: [65] [66] [67] , llm-line-node, text width=27.9em ] ] [ Adversarial Defense (§ 3.2), llm-line-node [ (i) Adversarial Detection: [68] [69](ii) Robust Inference: [70] , llm-line-node, text width=27.9em ] ] [ Jailbreak Attacks (§ 3.3), llm-line-node [ (i) Black-Box: [71] [72] [73] [74] [75] [76] [77] [78] [79] [80] [81] [82] [83] [84] [85] [86] [87] [88] [89]
@@
[90][91] [92] [93] [94] [95] , llm-line-node, text width=27.9em ] ] [ Jailbreak Defenses (§ 3.4), llm-line-node [ (i) Input Defense: [96] [97] [98] [99] [100](ii) Output Defense: [101] [102] [103]
@@
(iii) Robust Inference: [104] [105](iv) Ensemble Defense: [106] [107] , llm-line-node, text width=27.9em ] ] [ Prompt Injection Attacks (§ 3.5), llm-line-node [ (i) Black-Box: [108] [109] [110] [111] [112] [113] [114] [115] [116] [112] [113] [117] [118] , llm-line-node, text width=27.9em ] ] [ Prompt Injection Defenses (§ 3.6), llm-line-node [ (i) Input Defense: [119] [120](ii) Parameter Defense: [121] [122] [123] , llm-line-node, text width=27.9em ] ] [ Backdoor Attacks (§ 3.7), llm-line-node [ (i) Data Poisoning: [124] [125] [126] [127] [128] [129] [130] [131] [132] [133] [134] [135] [136] [137] [138]
@@
(ii) Training Manipulation: [139] [140] [141] [142](iii) Parameter Modification: [143] , llm-line-node, text width=27.9em ] ] [ Backdoor Defenses (§ 3.8), llm-line-node [ (i) Backdoor Detection: [144] [145] [146] [147] [148](ii) Backdoor Removal: [149] [150] [151] [152]
@@
(iii) Robust Training: [153] [154] [155](iv) Robust Inference: [156] [157] , llm-line-node, text width=27.9em ] ] [ Safety Alignment (§ 3.9), llm-line-node [ (i) Human Feedback: [158] [159] [160] [161] [162, 163] [164] [165] [166]
@@
(ii) AI Feedback: [167] [168] [169](iii) Social Interactions: [170] [171] , llm-line-node, text width=27.9em ] ] [ Energy Latency Attacks (§ 3.10), llm-line-node [ (i) White-Box: [172] [173] [174] [175](ii) Black-Box: [176] [177] , llm-line-node, text width=27.9em ] ] [ Model Extraction Attacks (§ 3.11), llm-line-node [ (i) Fine-tuning Stage: [178] [179](ii) Alignment Stage: [180] , llm-line-node, text width=27.9em ] ] [ Data Extraction Attacks (§ 3.12), llm-line-node [ (i) Black-Box: [181] [182] [183] [184] [185] [186] [187] [188] [189] [190](ii) White-Box: [191] , llm-line-node, text width=27.9em ] ] ] [ Vision-Language Pre-Training Models (§ 4) , vlp-bg-node [ Adversarial Attacks (§ 4.1), vlp-line-node [ (i) White-Box: [192] [193] [194](ii) Black-Box: [195] [196] [197] [198] [199] [200] [201] [202] , vlp-line-node, text width=27.9em ] ] [ Adversarial Defenses (§ 4.2), vlp-line-node [ (i) Adversarial Tuning: [203] [204] [205] [206] [207] [208] [209] [210] [211] [212] [213] [214]
@@
(ii) Adversarial Training: [215] [216](iii) Adversarial Detection: [217] [218] , vlp-line-node, text width=27.9em ] ] [ Backdoor & Poisoning Attacks (§ 4.3), vlp-line-node [ (i) Backdoor: [219] [220] [221] [222](ii) Poisoning: [223](iii) Backdoor & Poisoning: [224] , vlp-line-node, text width=27.9em ] ] [ Backdoor & Poisoning Defenses (§ 4.4), vlp-line-node [ (i) Backdoor Removal: [225] [226](ii) Robust Training: [227](ii) Backdoor Detection: [228] [229] [230] [231] [232] , vlp-line-node, text width=27.9em ] ] ] [ Vision-Language Models (§ 5) , vlm-bg-node [ Adversarial Attacks (§ 5.1), vlm-line-node [ (i) White-Box: [233] [234] [235] [236] [237](ii) Gray-Box: [238](iii) Black-Box: [239] [240] [241] [242] , vlm-line-node, text width=27.9em ] ] [ Jailbreak Attacks (§ 5.2), vlm-line-node [ (i) White-Box: [243] [244] [245] [246] [247] [248](ii) Black-Box: [249] [250] [251] [252] [253] , vlm-line-node, text width=27.9em ] ] [ Jailbreak Defenses (§ 5.3), vlm-line-node [ (i) Jailbreak Detection: [254] [255](ii) Jailbreak Prevention: [256] [257] [258] [259] [260] , vlm-line-node, text width=27.9em ] ] [ Energy Latency Attacks (§ 5.4), vlm-line-node [ (i) White-Box: [261] , vlm-line-node, text width=27.9em ] ] [ Prompt Injection Attack (§ 5.5), vlm-line-node [ (i) White-Box: [262](ii) Black-Box: [263] , vlm-line-node, text width=27.9em ] ] [ Backdoor & Poisoning Attacks (§ 5.6), vlm-line-node [ (i) Backdoor: [264] [265] [266] [267](i) Poisoning: [268] , vlm-line-node, text width=27.9em ] ] ] [ Diffusion Models (§ 6) , dm-bg-node [ Adversarial Attacks (§ 6.1), dm-line-node [ (i) White-Box: [269] [270](ii) Gray-Box: [271] [272](iii) Black-Box: [273] [274] [275] [276] [277] [278] [279] , dm-line-node, text width=27.9em ] ] [ Jailbreak Attacks (§ 6.2), dm-line-node [ (i) White-Box: [280] [281] [282] [283](ii) Gray-Box: [284] [285](iii) Black-Box: [286] [287] [288] [289] [290] [291] , dm-line-node, text width=27.9em ] ] [ Jailbreak Defenses (§ 6.3), dm-line-node [ (i) Concept Erasure: [292] [293] [294] [295] [296] [297] [298] [299] [300] [301] [302] [303] [304] [305] [306] [307]
@@
[308] [309] [310] [311] [312] [313](ii) Inference Guidance: [314] [315] [316] , dm-line-node, text width=27.9em ] ] [ Backdoor Attacks (§ 6.4), dm-line-node [ (i) Training Manipulation: [317] [318] [319] [320] [321]
@@
(ii) Data Poisoning: [322] [323] [324] [325] [326] [327] [328] [329] , dm-line-node, text width=27.9em ] ] [ Backdoor Defenses (§ 6.5), dm-line-node [ (i) Backdoor Detection: [330] [331] [332](ii) Backdoor Removal: [333] [334] [335] [336] , dm-line-node, text width=27.9em ] ] [ Membership Inference Attacks (§ LABEL:sec:dm_membership_inference_attacks), dm-line-node [ (i) White-Box: [337] [338] [339] [340](ii) Gray-Box: [341] [342] [343] [344] [345] [346]
@@
(iii) Black/White-Box: [339](iii) Black-Box: [347] [348] [349] [350] [351] , dm-line-node, text width=27.9em ] ] [ Data Extraction Attacks (§ LABEL:sec:dm_data_extraction_attacks), dm-line-node [ (i) Explicit Condition-based Extraction: [352] [353](ii) Surrogate Condition-based Extraction: [354] [355] , dm-line-node, text width=27.9em ] ] [ Model Extraction Attacks (§ LABEL:sec:dm_model_extraction_attacks), dm-line-node [ (i) LoRA-Based Extraction: [356] , dm-line-node, text width=27.9em ] ] [ Intellectual Property Protection (§ LABEL:sec:dm_intellectual_property_protection), dm-line-node [ (i) Natural Data Protection: [357] [358] [359] [360] [361] [362] [363] [364] [365] [366] [367] [368] [369]
@@
(ii) Generated Data Protection: [370] [371] [372] [373](ii) Model Protection: [374] [375] [376] [377] [378][379] , dm-line-node, text width=27.9em ] ] ] [ Agent (§ LABEL:sec:agent) , agent-bg-node [ LLM Agent (§ LABEL:sec:agent_LLM), agent-line-node [ (i) Attacks: [380] [381] [382] [383] [384] [385](i) Defenses: [386] [387] [388](i) Benchmarks: [389] [390] [391] , agent-line-node, text width=27.9em ] ] [ VLM Agent (§ LABEL:sec:agent_VLM), agent-line-node [ (i) Attacks: [392] [393] [394] [395] , agent-line-node, text width=27.9em ] ] ] ]
Figure 3: Organization of this survey.
@@
Survey Year Model Topic Zhang et al. [396] 2024 VFM&VLP&VLM Adversarial, Jailbreak, Prompt Injection Truong et al. [397] 2024 DM Adversarial, Backdoor, Membership inference Zhao et al. [398] 2024 LLM Backdoor Yi et al. [399] 2024 LLM Jailbreak Jin et al. [400] 2024 LLM&VLM Jailbreak Liu et al. [401] 2024 VLM Jailbreak Liu et al. [402] 2024 VLM Adverarial, Backdoor, Jailbreak, Prompt Injection Cui et al. [403] 2024 LLM agent Adversarial, Backdoor, Jailbreak Deng et al. [404] 2024 LLM agent Adversarial, Backdoor, Jailbreak Gan et al. [405] 2024 LLM&VLM agent Adversarial, Backdoor, Jailbreak
Our survey methodology is structured as follows. First, we conducted a keyword-based search targeting specific model types and threat types to identify relevant papers. Next, we manually filtered out non-safety-related and non-technical papers. For each remaining paper, we categorized its proposed method or framework by analyzing its settings and attack/defense types, assigning them to appropriate categories and subcategories. In total, we collected 390 technical papers, with their distribution across years, model types, and attack/defense strategies illustrated in Figure 1. As shown, safety research on large models has surged significantly since 2023, following the release of ChatGPT. Among the model types, LLMs and DMs have garnered the most attention, accounting for over 60% of the surveyed papers. Regarding attack types, jailbreak, adversarial, and backdoor attacks were the most extensively studied. On the defense side, jailbreak defenses received the highest focus, followed by adversarial defenses. Figure 2 presents a cross-view of temporal trends across model types and attack/defense categories, offering a detailed breakdown of the reviewed works. Notably, research on attacks constitutes 60% of the studied. In terms of defense, while defense research accounts for only 40%, underscoring a significant gap that warrants increased attention toward defense strategies. The overall structure of this survey is outlined in Figure 3.
Difference to Existing Surveys. Large-model safety is a rapidly evolving field, and several surveys have been conducted to advance research in this area. Recently, Slattery et al. [406] introduced an AI risk framework with a systematic taxonomy covering all types of risks. In contrast, our focus is on the technical aspects, specifically the attack and defense techniques proposed in the literature. Table I lists the technical surveys we identified, each focusing on a specific type of model or threat (e.g., LLMs, VLMs, or jailbreak attacks/defenses). Compared with these works, our survey offers both a broader scope—covering a wider range of model types and threats—and a higher level perspective, centering on overarching methodologies rather than minute technical details.

### v5

@@ Abstract
The rapid advancement of large models, driven by their exceptional abilities in learning and generalization through large-scale pre-training, has reshaped the landscape of Artificial Intelligence (AI). These models are now foundational to a wide range of applications, including conversational AI, recommendation systems, autonomous driving, content generation, medical diagnostics, and scientific discovery. However, their widespread deployment also exposes them to significant safety risks, raising concerns about robustness, reliability, and ethical implications. This survey provides a systematic review of current safety research on large models, covering Vision Foundation Models (VFMs), Large Language Models (LLMs), Vision-Language Pre-training (VLP) models, Vision-Language Models (VLMs), Diffusion Models (DMs), and large-model-powered Agents. Our contributions are summarized as follows: (1) We present a comprehensive taxonomy of safety threats to these models, including adversarial attacks, data poisoning, backdoor attacks, jailbreak and prompt injection attacks, energy-latency attacks, data and model extraction attacks, and emerging agent-specific threats. (2) We review defense strategies proposed for each type of attacks if available and summarize the commonly used datasets and benchmarks for safety research. (3) Building on this, we identify and discuss the open challenges in large model safety, emphasizing the need for comprehensive safety evaluations, scalable and effective defense mechanisms, and sustainable data practices. More importantly, we highlight the necessity of collective efforts from the research community and international collaboration. Our work can serve as a useful reference for researchers and practitioners, fostering the ongoing development of comprehensive defense systems and platforms to safeguard AI models. GitHub: https://github.com/xingjunm/Awesome-Large-Model-Safety.
@@
AI Safety, Large Model Safety, Agent Safety, Attacks and Defenses
@@ 1 Introduction
Figure 1: Left: The number of surveyed technical papers on attacks, defenses, and benchmarks/datasets. Middle: Distribution of surveyed technical papers by model type. Right: Distribution of surveyed technical papers by attack and defense type.
Figure 2: Left: The quarterly trend in the number of surveyed safety papers across different models; Middle: Proportional distribution of attack and defense studies associated with large models. Right: Annual trend in the number of surveyed safety papers on various attacks and defenses, ordered from most to least studied.
@@
forked edges, for tree= grow=east, reversed=true, anchor=mid west, parent anchor=mid east, child anchor=mid west, base=left, font=, rectangle, draw=none, sharp corners, align=left, minimum width=2em, edge+=darkgray, line width=0.6pt, s sep=2.8pt, inner xsep=1.8pt, inner ysep=2.7pt, line width=0.7pt, par/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=0font=,, where level=1text width=4.3em,font=,, where level=2text width=9.7em,font=,, [ Large Model and Agent Safety, par [ Vision Foundation Models (§ 2) , vfm-bg-node [ Attacks and Defenses for ViT (§ 2.1), vfm-line-node [ (i) Adversarial Attacks: [1] [2] [3] [4] [5] [6] [7] [8] [9] [10] [11] [12] [13] [14] [15] [16] [17] [18] [19]
@@
(iii) Backdoor Attacks: [39] [40] [41] [42] [43](iv) Backdoor Defenses: [44] [45] [46] , vfm-line-node, text width=32.7em ] ] [ Attacks and Defenses for SAM (§ 2.2), vfm-line-node [ (i) Adversarial Attacks: [47] [48] [49] [50] [51] [52] [53] [54](ii) Adversarial Defenses: [55]
@@
(iii) Backdoor & Poisoning Attacks: [56] [57] , vfm-line-node, text width=32.7em ] ] ] [ Large Language Models (§ 3) , llm-bg-node [ Adversarial Attack (§ 3.1), llm-line-node [ (i) White-Box: [58] [59] [60] [61] [62] [63] [64](ii) Black-Box: [65] [66] [67] , llm-line-node, text width=32.7em ] ] [ Adversarial Defense (§ 3.2), llm-line-node [ (i) Adversarial Detection: [68] [69](ii) Robust Inference: [70] , llm-line-node, text width=32.7em ] ] [ Jailbreak Attacks (§ 3.3), llm-line-node [ (i) Black-Box: [71] [72] [73] [74] [75] [76] [77] [78] [79] [80] [81] [82] [81] [83] [84] [85] [86] [87] [88] [89] [90] [91]
@@
[92] [93] [94] [95] [96] [97] [98] (ii) White-Box: [99] [100] , llm-line-node, text width=32.7em ] ] [ Jailbreak Defenses (§ 3.4), llm-line-node [ (i) Input Defense: [101] [102] [103] [104] [105] [106] [107] [108] [109](ii) Output Defense: [110] [111] [112] [113] [114]
@@
(iii) Ensemble Defense: [115] [116] [117] [118] (iv) Defenses against Fine-tuning Attacks: [119] [120] [121] [122] [123] [124] , llm-line-node, text width=32.7em ] ] [ Prompt Injection Attacks (§ 3.5), llm-line-node [ (i) Black-Box: [125] [126] [127] [128] [129] [130] [131] [132] [133] [129] [130] [134] [135] [136] , llm-line-node, text width=32.7em ] ] [ Prompt Injection Defenses (§ 3.6), llm-line-node [ (i) Input Defense: [137] [138](ii) Parameter Defense: [139] [140] [141] , llm-line-node, text width=32.7em ] ] [ Backdoor Attacks (§ 3.7), llm-line-node [ (i) Data Poisoning: [142] [143] [144] [145] [146] [147] [148] [149] [150] [151] [152] [153] [154] [155] [156] [157]
@@
(ii) Training Manipulation: [158] [159] [160] [161](iii) Parameter Modification: [162] , llm-line-node, text width=32.7em ] ] [ Backdoor Defenses (§ 3.8), llm-line-node [ (i) Backdoor Detection: [163] [164] [165] [166] [167] [168](ii) Backdoor Removal: [169] [170] [171] [172]
@@
(iii) Robust Training: [173] [174] [175](iv) Robust Inference: [176] [177] [178] , llm-line-node, text width=32.7em ] ] [ Safety Alignment (§ 3.9), llm-line-node [ (i) Human Feedback: [179] [180] [181] [182] [183, 184] [185] [186] [187]
@@
(ii) AI Feedback: [188] [189] [190](iii) Social Interactions: [191] [192] (iv) Deceptive Alignment: [193] [194] [195] , llm-line-node, text width=32.7em ] ] [ Energy Latency Attacks (§ 3.10), llm-line-node [ (i) White-Box: [196] [197] [198] [199] [200](ii) Black-Box: [201] [202] , llm-line-node, text width=32.7em ] ] [ Model Extraction Attacks (§ 3.11), llm-line-node [ (i) Fine-tuning Stage: [203] [204](ii) Alignment Stage: [205] , llm-line-node, text width=32.7em ] ] [ Data Extraction Attacks (§ 3.12), llm-line-node [ (i) Black-Box: [206] [207] [208] [209] [210] [211] [212] [213] [214] [215] [216](ii) White-Box: [217] , llm-line-node, text width=32.7em ] ] ] [ Vision-Language Pre-Training Models (§ 4) , vlp-bg-node [ Adversarial Attacks (§ 4.1), vlp-line-node [ (i) White-Box: [218] [219] [220] [221](ii) Black-Box: [222] [223] [224] [225] [226] [227] [228] [229][230] [231] [232] , vlp-line-node, text width=32.7em ] ] [ Adversarial Defenses (§ 4.2), vlp-line-node [ (i) Adversarial Tuning: [233] [234] [235] [236] [237] [238] [239] [240] [241] [242] [243] [244]
@@
(ii) Adversarial Training: [245] [246](iii) Adversarial Detection: [247] [248] , vlp-line-node, text width=32.7em ] ] [ Backdoor & Poisoning Attacks (§ 4.3), vlp-line-node [ (i) Backdoor: [249] [250] [251] [252](ii) Poisoning: [253] [254](iii) Backdoor & Poisoning: [255] , vlp-line-node, text width=32.7em ] ] [ Backdoor & Poisoning Defenses (§ 4.4), vlp-line-node [ (i) Backdoor Removal: [256] [257](ii) Robust Training: [258](ii) Backdoor Detection: [259] [260] [261] [262] [263] , vlp-line-node, text width=32.7em ] ] ] [ Vision-Language Models (§ 5) , vlm-bg-node [ Adversarial Attacks (§ 5.1), vlm-line-node [ (i) White-Box: [264] [265] [266] [267] [268](ii) Gray-Box: [269](iii) Black-Box: [270] [271] [272] [273] , vlm-line-node, text width=32.7em ] ] [ Jailbreak Attacks (§ 5.2), vlm-line-node [ (i) White-Box: [274] [275] [276] [277] [278] [279](ii) Black-Box: [280] [281] [282] [283] [284] , vlm-line-node, text width=32.7em ] ] [ Jailbreak Defenses (§ 5.3), vlm-line-node [ (i) Jailbreak Detection: [285] [286](ii) Jailbreak Prevention: [287] [288] [289] [290] [291] , vlm-line-node, text width=32.7em ] ] [ Energy Latency Attacks (§ 5.4), vlm-line-node [ (i) White-Box: [292] , vlm-line-node, text width=32.7em ] ] [ Prompt Injection Attack (§ 5.5), vlm-line-node [ (i) White-Box: [293](ii) Black-Box: [294] , vlm-line-node, text width=32.7em ] ] [ Backdoor & Poisoning Attacks (§ 5.6), vlm-line-node [ (i) Backdoor: [295] [296] [297] [298](i) Poisoning: [299] , vlm-line-node, text width=32.7em ] ] ] [ Diffusion Models (§ 6) , dm-bg-node [ Adversarial Attacks (§ 6.1), dm-line-node [ (i) White-Box: [300] [301] [302](ii) Gray-Box: [303] [304] [305] [306]
(iii) Black-Box: [307] [308] [309] [310] [311] [312] [313] , dm-line-node, text width=32.7em ] ] [ Jailbreak Attacks (§ 6.2), dm-line-node [ (i) White-Box: [314] [315] [316] [317](ii) Gray-Box: [318] [319]
(iii) Black-Box: [320] [321] [322] [323] [324] [325] [326] [327] , dm-line-node, text width=32.7em ] ] [ Jailbreak Defenses (§ 6.3), dm-line-node [ (i) Concept Erasure: [328] [329] [330] [331] [332] [333] [334] [335] [336] [337] [338] [339] [340] [341] [342] [343] [344] [345]
@@
[346] [347] [348] [349] [350] [351](ii) Inference Guidance: [352] [353] [354] [355] [356] , dm-line-node, text width=32.7em ] ] [ Backdoor Attacks (§ 6.4), dm-line-node [ (i) Training Manipulation: [357] [358] [359] [360] [361]
@@
(ii) Data Poisoning: [362] [363] [364] [365] [366] [367] [368] [369] , dm-line-node, text width=32.7em ] ] [ Backdoor Defenses (§ 6.5), dm-line-node [ (i) Backdoor Detection: [370] [371] [372](ii) Backdoor Removal: [373] [374] [375] [376] [377] , dm-line-node, text width=32.7em ] ] [ Membership Inference Attacks (§ 6.6), dm-line-node [ (i) White-Box: [378] [379] [380] [381](ii) Gray-Box: [382] [383] [384] [385] [386] [387]
@@
(iii) Black/White-Box: [380](iii) Black-Box: [388] [389] [390] [391] [392] , dm-line-node, text width=32.7em ] ] [ Data Extraction Attacks (§ 6.7), dm-line-node [ (i) Explicit Condition-based Extraction: [393] [394](ii) Surrogate Condition-based Extraction: [395] [396] , dm-line-node, text width=32.7em ] ] [ Model Extraction Attacks (§ 6.8), dm-line-node [ (i) LoRA-Based Extraction: [397] , dm-line-node, text width=32.7em ] ] [ Intellectual Property Protection (§ 6.9), dm-line-node [ (i) Natural Data Protection: [398] [399] [400] [401] [402] [403] [404] [405] [406] [407] [408] [409] [410]
@@
(ii) Generated Data Protection: [411] [412] [413] [414](ii) Model Protection: [415] [416] [417] [418] [419][420] , dm-line-node, text width=32.7em ] ] ] [ Agent (§ 7) , agent-bg-node [ Indirect Prompt Injection (§ 7.17.2), agent-line-node [ (i) Malicious Instruction [127] [421] [422] [125] [126] [423](ii) Indirect Jailbreak [424] [425] [426]
(iii) Defenses [427] [428] [429] [430] [431] [432] [433] , agent-line-node, text width=32.7em ] ] [ Memory Attacks & Defenses (§ 7.3-7.4), agent-line-node [ (i) Backdoor Attacks: [434] [435] [436] [364] [437] [438] [439] [440](ii) Poisoning Attacks: [441] [442] [443] [444]
(iii) Defenses: [427] [445] [446] [447] [448] [449] [450] [451] , agent-line-node, text width=32.7em ] ] [ Tool Attacks & Defenses (§ 7.57.6), agent-line-node [ (i) Tool Manipulation: [452] [453] [454] [455] [456](ii) MCP Manipulation: [457] [458] [459]
(iii) Defenses: [460] [461] [462] [463] , agent-line-node, text width=32.7em ] ] [ VLM Agent (§ 7.7-7.8), agent-line-node [ (i) Attacks: [464] [465] [466] [467] [468] [469] [470] [471](ii) Defenses: [472] [473] [474] [475] , agent-line-node, text width=32.7em ] ] [ Multi-Agent Systems (§ 7.9-7.10), agent-line-node [ (i) Attacks: [476] [477] [478] [479] [480] [481] [482] [483] [484](ii) Defenses: [485] [486] [487] [488] [489] [484] , agent-line-node, text width=32.7em ] ] [ Embodied Agents (§ 7.11-7.12), agent-line-node [ (i) Attacks: [490] [491] [492] [493] [494] [495] [496] [497] [498] [499] [500] [501] [502] [503]
(ii) Defenses: [504] [505] [506] [507] , agent-line-node, text width=32.7em ] ] [ Agentic Attacks & Defenses (§ 7.13), agent-line-node [ (i) Agentic Attacks: [508] [509] [510] [511] [512] [513](ii) Agentic Defenses: [514] [515] [516] , agent-line-node, text width=32.7em ] ] [ Benchmarks (§ 7.14), agent-line-node [ (i) Simulation-based Benchmarks: [140] [517] [518] [519] [520] [521] [522] [523] [524] [525] [526] [527] [528]
(ii) Real Interaction Benchmarks: [529] [530] [531] [532] [533] [534] [535] [536] [537] [538] [539] [540] [541] [542] [543] [544] , agent-line-node, text width=32.7em ] ] ] ]
Figure 3: A road map of this survey.
@@
Survey Year VFM VLP LLM VLM DM LLM-Agent VLM-Agent Zhang et al. [545] 2024 Truong et al. [546] 2024 Zhao et al. [547] 2024 Yi et al. [548] 2024 Jin et al. [549] 2024 Liu et al. [550] 2024 Liu et al. [551] 2024 Cui et al. [552] 2024 Gan et al. [553] 2024 Deng et al. [554] 2025 Ye et al. [555] 2025 Wang et al. [556] 2025 Our Survey 2025
Our survey methodology is structured as follows. First, we conducted a keyword-based search targeting specific model types and threat types to identify relevant papers. Next, we manually filtered out non-safety-related and non-technical papers. For each remaining paper, we categorized its proposed method or framework by analyzing its settings and attack/defense types, assigning them to appropriate categories and subcategories. In total, we reviewed 574 technical papers, with their distribution across years, model types, and attack/defense strategies illustrated in Figure 1. As shown, safety research on large models has surged significantly since 2023, following the release of ChatGPT. Among the model types, LLMs, DMs and Agents have garnered the most attention, accounting for 71.32% of the surveyed papers. Regarding attack types, jailbreak, adversarial, and backdoor attacks were the most extensively studied. On the defense side, jailbreak defenses received the highest focus, followed by adversarial defenses. Figure 2 presents a cross-view of temporal trends across model types and attack/defense categories, offering a detailed breakdown of the reviewed works. Notably, research on attacks constitutes $\sim$60of the studied. In terms of defense, while defense research accounts for only $\sim$40%, underscoring a significant gap that warrants increased attention toward defense strategies. The overall structure of this survey is outlined in Figure 3.
Difference to Existing Surveys. Large-model safety is a rapidly evolving field, and several surveys have been conducted to advance research in this area. Recently, Slattery et al. [557] introduced an AI risk framework with a systematic taxonomy covering all types of risks. In contrast, our focus is on the technical aspects, specifically the attack and defense techniques proposed in the literature. Table I summarizes the technical surveys we identified, each concentrating on a few model types or threat categories (e.g., LLMs, VLMs, agents, or jailbreak attacks/defenses). Compared with these works, our survey provides both a broader scope by covering a wider range of model types and threats, and a more high-level perspective that focuses on overarching methodologies rather than specific technical details.